OpenStudy (cicilybailey14):

Help! Find the slope of the line passing through the pair of points. If the line is horizontal or vertical, state that. (-2, 5), (8, -4)

OpenStudy (one098):

Do you know how to find slope?

OpenStudy (one098):

Use m for slope: \[m=\frac{ rise }{ run }\]

OpenStudy (cicilybailey14):

okay...

OpenStudy (igreen):

Use the slope formula.. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug in the points: \(m = \dfrac{-4-5}{8+2}\) Simplify that.

OpenStudy (cicilybailey14):

-0.8 @iGreen

OpenStudy (cicilybailey14):

waittt

OpenStudy (igreen):

Actually you made a mistake..

OpenStudy (cicilybailey14):

oops hold on

OpenStudy (cicilybailey14):

-0.9?

OpenStudy (igreen):

Yes, that's your slope.. -0.9 or \(-\dfrac{9}{10}\)